40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 22 02:40:12 2013, Jenne, Update, LSC, DRMI Locked for 1+ minute!!!!!! DRMI_First_Lock_1plusMinute_21Aug2013.png
    Reply  Thu Aug 22 07:57:57 2013, Lisa, Update, LSC, DRMI Locked for 1+ minute!!!!!! 
       Reply  Thu Aug 22 13:03:40 2013, Jenne, Update, LSC, DRMI Locked for 1+ minute!!!!!! 
    Reply  Thu Aug 22 10:20:32 2013, kiwamu, Update, LSC, DRMI Locked for 1+ minute!!!!!! 
    Reply  Thu Aug 22 13:20:54 2013, Koji, Update, LSC, DRMI Locked for 1+ minute!!!!!! 
    Reply  Sat Aug 24 00:11:07 2013, Koji, Update, LSC, DRMI Locked with improved lock streatches Screenshot-Untitled_Window.png
       Reply  Sun Aug 25 06:07:11 2013, rana, Update, LSC, DRMI Locked with improved lock streatches 
Message ID: 9049     Entry time: Thu Aug 22 02:40:12 2013     Reply to this: 9050   9051   9053   9060
Author: Jenne 
Type: Update 
Category: LSC 
Subject: DRMI Locked for 1+ minute!!!!!! 

[Jenne, Koji]

The DRMI has been locked!!  And at least one time, it was for more than one minute!! 

DRMI_First_Lock_1plusMinute_21Aug2013.png

We are not 100% sure yet that it's correctly sideband locked.  The test of this was to put a 50% BS in front of the AS camera (so after the beam has gone to AS55), and send the light over to a PDA10CF Thorlabs PD.  I locked the Michelson on carrier for the alignment of this diode.  Then I strung a cable to the control room, and plugged it into the RF spectrum analyzer.  (First, I had turned off the green beat PD power, so there wasn't any RF stuff on the line that I unplugged).  It's hard to watch the screen and a tv / dataviewer at the same time, so I've taken a video, so that we can see the nicely locked round DRMI beam on the AS camera, and the spectrum analyzer.  My phone is working very hard at uploading the video, but we may have to wait until tomorrow for that.  However, I think that we're locked on the 55MHz sideband. (Also, maybe I'm too tired or excited or something, but how do you make the real cameras take video??)

EDIT:  Video uploaded.  Pause the video at 10 seconds, and you'll see that we've got a strong 110MHz peak!!  Hoooray!  The TV in the upper right side of the video is AS.  You can see as we flash, the peaks go up and down.  When there's no resonance, the 110 peak goes away. (Ex., when I'm PRMI locked on the sideband, there isn't a visible peak).

 

Alignment procedure was as normal:  Lock and align the arms. Misalign ETMs. Check that MICH fringes look good (ASS does a nice enough job that I don't actually lock and align the Michelson anymore).  Restore the PRM.  Lock PRMI.  Tweak PRM alignment to maximize POP110I.  At this point, Koji and I played a little with the PRMI, but when we finished with that, we restored the SRM, and tweaked its alignment by making nice overlap on the AS camera.

Then, we tried some DRMI settings, started seeing some locks, and played a bit with trying to optmize the settings that we have. 

DETAILS:

PRMI settings: 

PRCL ASC is on (with loop triggering).  MICH gain = -0.8, PRCL gain = +0.05.  FM4, FM5 always on, FM2 triggered.  Loop and filter module triggering on POP22I.  No power normalization.  MICH and PRCL locked on REFL55 I&Q, with 1's in the LSC input matrix.  PRCL actuating on PRM with +1, MICH actuating on BS with +0.5, PRM with -0.267. 

I took transfer functions between REFL55 I&Q and REFL11 I&Q, to determine the relative gains and signs.  REFL11I's gain should be -18dB relative to REFL55I, with the opposite sign.  We tried PRMI locking with MICH = 1*REFL55Q and PRCL = -0.125*REFL11I for the input matrix.  Still no power normalization (we haven't used power norm at all today, so I'll quit writing that).

I took transfer functions between REFL55 I&Q and REFL33 I&Q.  REFL33I's gain is -8dB relative to REFL55I, but they have the same sign.  We tried locking PRMI with MICH = 1*REFL55Q and PRCL = +0.6*REFL33I.  Success.

Next up, some Optickle simulations, to help us go in the right direction for DRMI locking.  I checked the signs of the error signals REFL55I (PRM sweep), REFL11I (PRM sweep) and REFL55Q (MICH sweep) in both PRMI and DRMI configurations.  For all of these cases, the signs were the same (i.e. no sign flips needed to happen for DRMI locking, relative to PRMI locking).  I checked the sensing matrices for DRMI and PRMI for those same signals, and took the ratios of the sensing matrix elements.  This gave me the ratio of optical gains for each error signal, in the DRMI case vs. PRMI case, so any servo gain changes should be the inverse of these numbers.  These numbers are all DRMI/PRMI:  REFL55I PRCL response = 0.76, REFL11I PRCL response = 0.99, REFL55Q MICH response = 18.  So, when trying to lock the DRMI, we wanted to keep the gains for PRCL about the same, reduce the servo gain for MICH by a factor of ~20, but keep the same signs for everything.

In doing that, we started seeing some short DRMI locks, so we twiddled some parameters (mostly the elements in the LSC input matrix) a bit.  We eventually settled on:  PRCL = -0.125*REFL11I, MICH = 0.1*REFL55Q, and SRCL = 1.0 * REFL55I.  The output matrix was the same (MICH pushing on BS and PRM, PRCL on PRM), with the addition of a +1 in the SRCL -> SRM element.  For all 3 degrees of freedom (PRCL, MICH, SRCL), FMs 4 and 5 were always on.  For PRCL, FMs 2,3,6 were triggered to come on after 0.5 seconds of delay.  The PRCL FM triggers helped enormously.  I tried several other things, including changing the MICH input matrix element up and down in value, changing the SRCL input matrix element up and down in value, and engaging triggering for a few different filters in the MICH and SRCL degrees of freedom.  However, none of these made things better, and several made things worse.  Most notably, for SRCL, engaging triggering for FMs 2 and 3 kicked the cavities out of lock, which implies that perhaps our gain isn't high enough yet (and thus our UGF isn't very high yet).  I changed FM1 of SRCL to be +3dB of gain (from +10dB), and it would live through that coming on (trigger delayed by 1 sec, then ramping up over 1 second), but within a second after the filter finishing coming on, the cavity would fall out of lock (not violently kicked, just not locked anymore). 

At this point, we were trying to figure out a way to confirm what kind of lock we had.  I checked Optickle again, and we do not expect to see a significant change in POP110I between the PRMI and DRMI cases, so that isn't a useful check.  We dreamed of having our AS110 demod board, or the AS OSA set up, but neither of those was going to happen tonight.  Instead, Koji suggested hooking up the PD, and looking directly at the output. 

To-do:  Set up the AS OSA.  Also, perhaps temporarily borrow the 110 demod board from POP.  We were triggering on POP22 tonight, and that seemed to work okay.

 

 

ELOG V3.1.3-